FM SPECTROSCOPY OF THE WATER VAPOR AT 1.3 μm

ANDREW BEZANT, \underline{PATRICK DUPRÉ}, and TERRY A. MILLER, Department of Chemistry, Ohio State University, 120 W 18^{th} Aue, Columbus OH, 43210.

In our laboratory we are particularly interested in developping sensitive spectroscopic techniques for the aim of study of reactive species. Moreover, the possibility of quantitative concentration of intermediates is a major key of understanding reactions mechanisms. For this purpose we set up a frequency-modulation absorption spectrometer based on a diode laser source working at 1.3 μm. Our first results concern the water molecule whose some known transitions (polyad 2 ν) are used to qualify the apparatus in term of linear response, sensitivity and absolute concentration by thermally controlling the water vapor pressure.